Let X, Y, Z have joint probability distribution function





f(x , y , z) = { k(x^2+ yz), 0 ≤ x ≤ 1, 0 ≤ y ≤ 1, 0 ≤ z ≤ 1



0, elsewhere





Find



(i) the value of k.



(ii) the conditional expectation of Z given X= x,Y= y

1
Expert's answer
2022-05-11T17:44:44-0400

The following condition must be satisfied: "\int_0^1\int_0^1\int_0^1f(x,y,z)dxdydz=1".

Compute the left side of the latter: "\int_0^1\int_0^1\int_0^1f(x,y,z)dxdydz=\int_0^1\int_0^1\int_0^1k(x^2+yz)dxdydz=k\int_0^1\int_0^1(\frac13x^3+xyz)|_0^1dydz=k\int_0^1\int_0^1(\frac13+yz)dydz=k\int_0^1|(\frac13y+\frac12y^2z)|_0^1dz=k\int_0^1(\frac13+\frac12z)dz=k(\frac13z+\frac14z^2)|_0^1=k(\frac13+\frac14)=\frac{7}{12}k."

(i) From the latter we find that: "k=\frac{12}{7}".

(ii) The aim is to find: "E(Z|X=x,Y=y)". Using the definition of conditional expectation we get: "E(Z|X=x,Y=y)=\int_{0}^{1}zf_{Z|X,Y}(z|x,y)dz=\frac{1}{f_{X,Y}(x,y)}\int_{0}^{1}zf(z,x,y)dz".

"f_{X,Y}(x,y)=\int_{0}^{1}k(x^2+yz)dz=k(x^2z+\frac12yz^2)|_0^1=k(x^2+\frac12y)."

"\int_{0}^{1}zf(z,x,y)dz=\int_{0}^{1}k(x^2z+yz^2)dz=k(\frac12x^2+\frac13y)"

We get: "E(Z|X=x,Y=y)=\frac{\frac12x^2+\frac13y}{x^2+\frac12y}".

Answer: (i) "k=\frac{12}{7}" (ii) "E(Z|X=x,Y=y)=\frac{\frac12x^2+\frac13y}{x^2+\frac12y}"